Question: cubed root of x squared=9
After cubing both sides, the 9 becomes 729
leaving the equation x squared=729
Please help solve for x
Any help appreciated!!
Thank you

Answer by stanbon(75887)

cube root of x squared=9
x^(2/3)=9
Raise both sides to the (3/2) power to get:
x=9^(3/2)
x=[9^(1/2)]^3
x=3^3=27
Cheers,
Stan H.